A f:X→Y where X,Y ∈ Z+ , and  0≤ X ≤23 and, 0≤Y≤59, X is hour of clock and Y is minute of digital clock, Proof the following

                                                       a.     Is f:X→Y is well defined

b.     Is f:X→Y a one-to-one correspondence 


1
Expert's answer
2021-02-24T07:49:44-0500

Let "f:X\to Y" where "x, y\in \mathbb Z^+", and "0\le x\le 23, 0\le y\le 59", "x" is hour of clock and "y" is minute of digital clock.


a.     Let us show that "f:X\u2192Y" is not well defined. Indead, if we consider "x=1" and "y_1=10,\ y_2=20", that is the hours "1:10,\ 1:20", we have that "x=1" maps to at least two values "y=10" and "y=20." Therefore, "f:X\u2192Y" is not well defined.


b.   Let us show that "f:X\u2192Y" is not a one-to-one correspondence. Indeed, if we consider "x_1=1" and "x_2=2,\ y_1=y_2=20", that is the hours "1:20,\ 2:20", we have that "x_1=1" and "x_2=2" map to the same value "y=20." Therefore, "f:X\u2192Y" is not a one-to-one correspondence.